• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings/R. Zielinski

    Log entry time 09:14:27 on March 09, 2012

    Entry number 365063

    This entry is a followup to: 364291

    Followups:

    keyword=T2 rate problem

    The rate on T2 was still not as expected (<1Hz for cosmics).  Since we
    had adjusted the of s2m/s1 recently, but hadn't adjusted the delay of the
    RT signal, we suspected that the strobe signal was now coming too early.
     By adjusting the delay of the strobe we were able to get a T2 rate of
    ~4Hz for cosmics.  We will continue to monitor it when beam comes back.
    


    A copy of this log entry has been emailed to: melissac, rbziel, vasulk